Calcium carbonate is a common ingredient in antacids that reduces the discomfort associated with acidic stomach or heartburn. Stomach acid is hydrocholoric acid, HCl.HCl.

What volume in milliliters (mL) of an HClHCl solution with a pH of 1.581.58 can be neutralized by 20.0 mg20.0 mg of CaCO3?

If the stomach contains 24.3 mL24.3 mL of pH 1.58pH 1.58 solution, will all of the acid be neutralized?

yes

no

What percentage of the acid is neutralized? If all of the acid is neutralized enter 100%.
Answer #1

Answer:-

This question is solved by using the simple concept of stoichiometry. Firstly a balanced chemical equation is written then volume is calculated.

The answer is given in the image,

Anjwers PH=1.58 pu=-log[4] CH+] = 10-P1 =10-1.58 (H+] = 0.0263 M-Thie] 240 + Caloz + cachy +CO2 +H2O los So mans of calo3 = 2